kekeman:

A lamppost, CAB, bent at point A after a storm. The tip of the lamppost touched the ground at point C, as shown below: Triangle ABC has measure of angle C equal to 45 degrees, measure of angle ABC equal to 90 degrees, and length of BC equal to 12 feet. What is the height, in feet, of the portion AB of the lamppost? 12 tan 45° 12 divided by tan 45 degrees 12 divided by cos 45 degrees 12 cos 45°

Ashely:

it says "as shown below" is there a ss for this

kekeman:

Here: https://snipboard.io/bW2ACF.jpg

AZ:

 We have an angle. We have the side opposite it and the side adjacent to it. Do we use sin, cos or tan? Remember sin \(\theta\) = opposite/ hypotenuse cos \(\theta\) = adjacent/ hypotenuse tan \(\theta\) = opposite/ adjacent

kekeman:

I am pretty sure we have to use "tan "

kekeman:

So i think the answer is the first one (which is 12 tan 45°)

jhonyy9:

so i say bc. an angle has measure 90 degree and the other one has 45 degree so from this result that the 3rd angle has a measure of 45 degree too what mean that this is an isoscele triangle do you know the properties of an isoscele triangle ?

kekeman:

Nah

ItzAgar:

@kekeman wrote:
Nah
you just asked for help and he explained it.

kekeman:

Ik that and he asked me if i knew the properties of an isosceles triangle and i said nah

kekeman:

@kekeman wrote:
So i think the answer is the first one (which is 12 tan 45°)
But i was saying that i feel like the answer would be

TurtlePrincess:

It always helps to draw out your triangle and label everything you know. Also remember the acronym Soh-Cah-Toa. Sin = opposite / hypotenuse Cos = adjacent / hypotenuse Tan = opposite / adjacent here are the 2 givens, 45 degrees and adjacent side 12, we are looking for the opposite side of the angle which means we use Tan

TurtlePrincess:

\[\tan(45)=opposite/12 \] now to get opposite by its self we multiply 12 to both sides

kekeman:

So with saying this the answer is the first one right

TurtlePrincess:

Correct :)

AZ:

@kekeman wrote:
So i think the answer is the first one (which is 12 tan 45°)
Yup you had nailed it! Good job :)

AZ:

What Jhonyy was trying to say was that since it's a 45-45-90 triangle The other side has to be 45 because of the ratios of a 45-45-90 triangle. Also tan 45 = 1 which means 12 tan 45 = 12 However, if you just used the trig functions like we did that was more than enough :)

1 attachment
kekeman:

Ohhhh ok i get it